# Cohomology ring of non-compact abelian arrangements

### v1

@@ 5. Non-unimodular case
$0=$ $\sum_{q\in\pi^{-1}(p)}\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)-\sum_{q\in\pi^{-1}(p)}\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)$
@@
Example 5.8.
{tikzpicture}{tikzpicture}
Figure 4. A picture of the non unimodular arrangement in Example 5.8 when $(a,b)=(1,0)$ on the left, and $(a,b)=(0,1)$ on the right.
@@
(see Figure 4). The generalized circuits are $C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\}$ and $X=C\sqcup F=\{1,2,3\}\sqcup\{4\}$.
@@
Notice that in the case $a=1$ the unimodular covering of $\mathcal{A}$ is the arrangement in Example 4.4, and the pullback of eq. (23) is the sum of equations (12) and (14).
Now consider the generalized circuit $X$, we get $2^{a}$ relations because the intersection $H_{1}\cap H_{2}\cap H_{3}\cap H_{4}=W\cap H_{4}$ has $2^{a}$ distinct connected components, $p_{1},\cdots,p_{2^{a}}$ and for any $i=1,\cdots,2^{a}$:
@@
Lemma 5.9.
The classes $\omega_{W,A}$ for all independent sets $A\subseteq E$ and all connected components $W$ of $\bigcap_{i\in A}H_{i}$, generate $H^{*}(M(\mathcal{A}))$ as a $H^{*}(G^{r})$-module.
Proof.
Fix $i\in E$, we proceed by induction using the exact sequence (of $\mathbb{Z}$-module) of Corollary 4.2 given by deletion and contraction with respect to $H_{i}$:
$0\to H^{*}(M(\mathcal{A}^{\prime}))\xrightarrow{j^{*}}H^{*}(M(\mathcal{A}))\xrightarrow{\operatorname{res}}H^{*-(a+b-1)}(M(\mathcal{A}^{\prime\prime}))\to 0$
Let $\mathcal{K}\subseteq H^{*}(M(\mathcal{A}))$ be the $H^{*}(G^{r})$-submodule generated by all $\omega_{W,A}$, $\mathcal{K}^{\prime}\subseteq H^{*}(M(\mathcal{A}^{\prime}))$ the $H^{*}(G^{r})$-submodule generated by all $\omega_{W,A}^{\prime}$ with $i\not\in A$, and $\mathcal{K}^{\prime\prime}\subseteq H^{*}(M(\mathcal{A}^{\prime\prime}))$ the $H^{*}(G^{r-1})$-submodule generated by all $\omega_{W,A}^{\prime\prime}$ with $i\in A$.
By definition of the classes $\omega_{W,A}$ we have $j_{*}(\omega_{W,A}^{\prime})=\omega_{W,A}$ and
$\operatorname{res}(\omega_{W,A})=\begin{cases}\omega^{\prime\prime}_{W,A\setminus\{i\}}&\textnormal{if }i\in A,\\ 0&\textnormal{otherwise.}\end{cases}$
Since the pullback map in cohomology $H^{*}(G^{r})\twoheadrightarrow H^{*}(G^{r-1})$ is surjective and the residue map is a morphism of $H^{*}(G^{r})$-modules, we have
{tikzcd}
By diagram chasing the inductive step follows, this completes the proof. ∎
@@
Let $R$ be the free $H^{*}(G^{r})$-module generated by the classes $\omega_{W,A}\in H^{*}(M(\mathcal{A}))$ with $A\subseteq E$ independent and $W$ connected component of $\bigcap_{i\in A}H_{i}$. We endow $R$ with a ring structure by defining the multiplication
@@
Notice that, for any $A,B\subseteq E$, and for any connected component $W$ of $\bigcap_{i\in A}H_{i}$, $\eta_{W,A,B}\in R$.
@@
  • Whenever $A\sqcup B$ and $A^{\prime}\sqcup B^{\prime}$ are not disjoint or $A\sqcup B\sqcup A^{\prime}\sqcup B^{\prime}$ is dependent,
    $\eta_{W,A,B}\eta_{W^{\prime},A^{\prime},B^{\prime}}=0$ (26)
  • For any generalized circuit $X=C\sqcup F$ with $C=C^{+}\sqcup C^{-}$, and for any connected component $Y$ of $\bigcap_{i\in X}H_{i}$,
@@
Proof.
The cohomology ring is a quotient of $R$ by Lemma 5.9. Relation (26) holds by Lemma 5.5 and relations (27), (28) by Lemma 5.7.
Let $I$ be the ideal given by relations (26)-(28). There is a surjective map $R/I\rightarrow H^{*}(M(\mathcal{A}))$. We prove that this map is an isomorphism by induction on the cardinality of $\mathcal{A}$. The base case is trivial.
Let us fix $i\in E$. Let $R^{\prime}$, $R^{\prime\prime}$, $I^{\prime}$ and $I^{\prime\prime}$ be defined analogously with respect to the deletion $\mathcal{A}^{\prime}$ and the contraction $\mathcal{A}^{\prime\prime}$ respectively. By induction hypothesis $R^{\prime}/I^{\prime}\cong H^{*}(M(\mathcal{A}^{\prime}))$ and $R^{\prime\prime}/I^{\prime\prime}\cong H^{*}(M(\mathcal{A}^{\prime\prime}))$. Let us consider the following diagram:
{tikzcd}
where:
    the first horizontal map is defined by the inclusions $R^{\prime}\subseteq R$ and $I^{\prime}\subseteq I$;
    the second horizontal map is induced by the map $g\colon R\rightarrow R^{\prime\prime}$ given by
    $\omega_{W,A}z\mapsto\begin{cases}\omega_{W,A\setminus i}\bar{z},&\text{ if }i\in A,\\ 0&\text{ otherwise,}\end{cases}$
    for any $z\in H^{*}(G^{r})$, where $\bar{z}$ is the image of $z$ through $r\colon H^{*}(G^{r})\rightarrow H^{*}(H_{i})$. The map is well defined since the image of $I$ is $I^{\prime\prime}$;
    the second row is the deletion-contraction exact sequence of Corollary 4.2.
The left square of diagram commutes because $j^{*}(\eta^{\prime}_{W,A,B})=\eta_{W,A,B}$, as the right one, since $\operatorname{res}(\eta_{W,A,B})=\eta_{W,A\setminus\{i\},B}^{{}^{\prime\prime}}$ if $i\in A$ and $\operatorname{res}(\eta_{W,A,B})=0$ otherwise. This implies that the first horizontal map is injective and the second one surjective. We now want to prove that the central vertical map is injective.
Firstly, observe that for any relation $r$ of type (27) and any generator $\eta_{W,A,B}$ the element $r\eta_{W,A,B}$ is linear combination of relations of types (26) and (27). In particular, the ideal $I$ is generated by (26) and (27) as $\mathbb{Z}$-module. It follows that the map $g_{|I}\colon I\twoheadrightarrow I^{\prime\prime}$ is surjective.
@@ 6. Configuration spaces
In this section, we apply the main result of this article to the cohomology algebra of ordered configuration spaces in $\mathbb{R}^{b}\times(S^{1})^{a}$.
@@
Now, we write down the presentation of $H^{*}(\operatorname{Conf}_{n}(G);\mathbb{Z})$ by using Theorem 5.11. The ground set of the associated matroid is the set $E=\{ij|i,j\in[n],i<j\}$ with the lexicographic order. The poset of layers is $\mathcal{L}(\mathcal{A}_{n})=\Pi_{[n]}$ the poset of partitions of $[n]$. The circuits of the arrangement are of the form $\{ij,ik,jk\}$ with $i,j,k\in[n]$ and $i<j<k$.
@@
Recall that in $H^{*}(G^{n})$, $\psi_{ik}=\psi_{ij}+\psi_{jk}$ and notice that relation (26) $\omega_{ij}\omega_{jk}\omega_{ik}=0$ follows by multiplying equation (30) (resp. eq. (31)) by $\omega_{ik}$. In the case $b=1$, correction terms appear in our formulas. In the article [CT78] and in this one, the element $\omega_{ij}$ represent the cohomological class of $H_{ij}$ and hence there is no canonical morphism between the two algebras $\operatorname{gr}_{F_{*}}H^{*}(\operatorname{Conf}_{n}(G;\mathbb{K}))$ and $H^{*}(\operatorname{Conf}_{n}(G);\mathbb{K})$.

### v2

@@ 5. Non-unimodular case
0 0=0 = $\sum_{q\in\pi^{-1}(p)}\sum_{\begin{subarray}{c}K\subseteq C^{-}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)-\sum_{q\in\pi^{-1}(p)}\sum_{\begin{subarray}{c}K\subseteq C^{+}\\ K\neq\emptyset\end{subarray}}(-1)^{|K|}c_{i_{K}}^{d}\eta^{U}_{X\setminus K,K\setminus i_{K}}(q)$
@@
Example 3.
[…]
@@
(see Figure 4). The sets of nullity 111 are $C=C^{+}\sqcup C^{-}=\{1,2\}\sqcup\{3\}$ and $X=C\sqcup F=\{1,2,3\}\sqcup\{4\}$.
@@
Notice that in the case $a=1$ the unimodular covering of $\mathcal{A}$ is the arrangement in Example 2, and the pullback of eq. (19) is the sum of equations (8) and (10).
Now consider a set $X$ of nullity 111, we get $2^{a}$ relations because the intersection $H_{1}\cap H_{2}\cap H_{3}\cap H_{4}=W\cap H_{4}$ has $2^{a}$ distinct connected components, $p_{1},\cdots,p_{2^{a}}$ and for any $i=1,\cdots,2^{a}$:
@@
Let $R$ be the free $H^{*}(G^{r})$-module generated by the classes $\omega_{W,A}\in H^{*}(M(\mathcal{A}))$ with $A\subseteq E$ independent set and $W$ connected component of $\bigcap_{i\in A}H_{i}$. We endow $R$ with a ring structure by defining the multiplication
@@
Notice that, for any $A,B\subseteq E$ and for any connected component $W$ of $\bigcap_{i\in A}H_{i}$, the element $\eta_{W,A,B}\in R$ is defined as $\omega_{W,A}\psi_{B}$.
@@
  • For any independent set $A\subseteq E$ and any connect component $W$ of $\cap_{i\in A}H_{i}$ and any $z\in\ker\left(H^{*}(G^{r};\mathbb{Z})\to H^{*}(W;\mathbb{Z})\right)$,
    $\omega_{W,A}z=0$ (22)
  • For any central set $X=C\sqcup F$ of nullity 111 with associated signed circuit $C=C^{+}\sqcup C^{-}$, and for any connected component $Y$ of $\bigcap_{i\in X}H_{i}$,
@@
Remark 6.
Its easy to see that eq. (22) implies $\omega_{W,A}\psi_{B}^{j}=0$ for all $j=1,\dots,a$ and all $B\subseteq E$ dependent on $A$. The viceversa, for essential arrangements, holds only under the assumption that $\mathcal{A}$ is unimodular or after tensoring with $\mathbb{Q}$.
of Theorem 5.9.
Consider the map $p\colon R\to H^{*}(M(\mathcal{A});\mathbb{Z})$ given by the pullback of $M(\mathcal{A})\to G^{r}$ and that sends the formal generators $\omega_{W,A}$ to the cohomology classes $\omega_{W,A}$. The map $p$ is a ring homomorphism because the product in $R$ (defined in eq. (21)) coincides with the product in cohomology (see Lemma 5.6).
@@
We prove that $p$ induces an isomorphism $R/I\to H^{*}(M(\mathcal{A});\mathbb{Z})$ by induction on the cardinality of $\mathcal{A}$. The base case $\mathcal{A}=\emptyset$ is trivial.
Let $S=R/J$ and fix $i\in E$. Let $S^{\prime}$, $S^{\prime\prime}$, $I^{\prime}$, $I^{\prime\prime}$ be the analog objects with respect to the deletion $\mathcal{A}^{\prime}$ and the contraction $\mathcal{A}^{\prime\prime}$ of the subvariety $H_{i}$. Consider the short exact sequence
[…]
given by the inclusion and the residue map $g$ defined by
$g(\omega_{W,A}z)=\begin{cases}\omega_{W,A\setminus i}\bar{z},&\text{ if }i\in A,\\ 0&\text{ otherwise,}\end{cases}$
[…]
With an abuse of notation, we call $I$ the image of the ideal $I$ in the quotient ring $S$. Consider the diagram
[…]
By induction hypothesis the first and last columns are exact. The second row is exact by the previous considerations and the last one by Corollary 4.2. The left square of diagram commutes because $j^{*}(\eta^{\prime}_{W,A,B})=\eta_{W,A,B}$, as the right one, since $\operatorname{res}(\eta_{W,A,B})=\eta_{W,A\setminus\{i\},B}^{{}^{\prime\prime}}$ if $i\in A$ and $\operatorname{res}(\eta_{W,A,B})=0$ otherwise.
The Snake Lemma implies that the map $p$ is surjective, and that
$0\to I^{\prime}\to\ker p\to I^{\prime\prime}\to 0$
is exact.
Now we prove $\ker p=I$, we already noticed that $\ker p\supseteq I$. Regarding the other inclusion, observe that for any relation $r$ of type (23) and any generator $\eta_{W,A,B}$ the element $r\eta_{W,A,B}$ is linear combination of relations of types (22) and (23). In particular, the ideal $I$ is generated by (22) and (23) as $\mathbb{Z}$-module. It follows that the map $g_{|I}\colon I\twoheadrightarrow I^{\prime\prime}$ is surjective. Moreover, $I^{\prime}\subseteq I$ and this implies $\ker p\subseteq I$.
This completes the inductive step and concludes the proof. ∎
6. Relation with [CDD+20]
Our presentation, specialized to toric arrangements, is similar to the one obtained in [CDD+20]. The integral cohomology ring was studied in [CDD+20, Section 7] without providing an explicit presentation, but showing that the cohomology is generated by the differential forms $\frac{m(A)}{m(A\cup B)}\eta_{W,A,B}$. Those cohomological classes $\eta_{W,A,B}$ correspond to our classes $\eta_{W,A,B}$; recall that they depend on the choice of an orientation of each hypertorus.
The rational cohomology was presented in [CDD+20, Theorem 6.13] using the differential forms $\overline{\eta}_{W,A,B}$. That differential forms are an average of $\eta_{W,A,B}$ over all possible orientation choice
$\overline{\eta}_{W,A,B}=\sum_{D\subseteq A}(-1)^{\lvert D\rvert}2^{\lvert A\setminus D\rvert}\frac{m(A\setminus D)}{m(A)}\eta_{W(A\setminus D),A\setminus D,B\cup D}.$ (26)
Therefore, their eq. (28) in [CDD+20, Theorem 6.13] corresponds to our multiplication rule in $R$ eq. (21) and to eq. (22), and their eq. (29) holds in $H^{*}(G^{r})\subset R$. The eq. (30) is equivalent to eq. (23) in a very intricate way that we are going to explain.
Consider $X$ a subset of nullity 111, $C=C^{+}\sqcup C^{-}$ be the oriented circuit in $X$ and $Y$ a connected component of $\cap_{i\in X}H_{i}$. We substitute eq. (26) in their eq. (30):
(30(30)( 30 ) $=\sum_{i\in C}\sum_{\begin{subarray}{c}B\subseteq C\setminus\{i\}\\ \lvert B\rvert\text{ even}\end{subarray}}(-1)^{\lvert A_{<i}\rvert+\lvert B\cap C^{-}\rvert}\frac{m(A)}{m(X\setminus\{i\})}\overline{\eta}_{W(A),A,B}$
$=\sum_{i\in C}\sum_{\begin{subarray}{c}B\subseteq C\setminus\{i\}\\ \lvert B\rvert\text{ even}\end{subarray}}\sum_{D\subseteq A}(-1)^{\lvert A_{<i}\rvert+\lvert B\cap C^{-}\rvert+\lvert D\rvert}2^{\lvert A\setminus D\rvert}\frac{m(A\setminus D)}{m(X\setminus\{i\})}\eta_{W(A\setminus D),A\setminus D,B\cup D},$
where $A=X\setminus(B\cup\{i\})$ and $W(A)$ is the unique connected component of $\cap_{i\in A}H_{i}$ containing $Y$.
Let $E=A\setminus D$ and notice that for any $i,j\in C$ we have
$c_{i}\frac{\eta_{W(E),E,X\setminus(E\cup\{i\})}}{m(X\setminus\{i\})}=c_{j}\frac{\eta_{W(E),E,X\setminus(E\cup\{j\})}}{m(X\setminus\{j\})}.$ (27)
Substituting eq. (27) we have
$(30)=\sum_{\begin{subarray}{c}E\subset X\end{subarray}}\sum_{\begin{subarray}{c}B\subseteq C\setminus E\\ \lvert B\rvert\text{ even}\end{subarray}}\sum_{i\in C\setminus(E\cup B)}(-1)^{\lvert A_{<i}\rvert+\lvert B\cap C^{-}\rvert+\lvert D\rvert}2^{\lvert E\rvert}\frac{m(E)}{m(X\setminus\{i\})}\eta_{W(E),E,X\setminus(E\cup\{i\})}$
$=\sum_{E\subset X}(-1)^{\lvert X\setminus E\rvert-1}2^{\lvert E\rvert}\frac{m(E)c_{j}\eta_{W(E),E,X\setminus(E\cup\{j\})}}{m(X\setminus\{j\})}\sum_{\begin{subarray}{c}B\subseteq C\setminus E\\ \lvert B\rvert\text{ even}\end{subarray}}\sum_{i\in C\setminus(E\cup B)}(-1)^{\lvert(E\sqcup D)_{<i}\rvert+\lvert B\cap C^{-}\rvert}c_{i},$
where $j$ is any element in $X\setminus E$. Notice that $c_{i}=(-1)^{\lvert X_{<i}\rvert+\lvert\{i\}\cap C^{-}\rvert}$ and we claim that
$\sum_{\begin{subarray}{c}B\subseteq C\setminus E\\ \lvert B\rvert\text{ even}\end{subarray}}\sum_{i\in C\setminus(E\cup B)}(-1)^{\lvert B_{<i}\rvert-1+\lvert(B\cup\{i\})\cap C^{-}\rvert}=\begin{cases}-2^{\lvert C\setminus E\rvert-1}&\text{if }\emptyset\neq C\setminus E\subseteq C^{+},\\ 2^{\lvert C\setminus E\rvert-1}&\text{if }\emptyset\neq C\setminus E\subseteq C^{-},\\ 0&\text{otherwise.}\end{cases}$ (28)
The left hand side of eq. (28) is equal to
$\sum_{\begin{subarray}{c}\tilde{B}\subseteq C\setminus E\\ \lvert\tilde{B}\rvert\text{ odd}\end{subarray}}(-1)^{\lvert\tilde{B}\cap C^{-}\rvert-1}$
where $\tilde{B}=B\sqcup\{i\}$, and so proving eq. (28) becomes an exercise left to the reader.
Finally, we have
$\begin{split}(30)=&\sum_{\begin{subarray}{c}E\subset X\\ \emptyset\neq C\setminus E\subseteq C^{-}\end{subarray}}(-1)^{\lvert X\setminus E\rvert-1}2^{\lvert E\cup C\rvert-1}\frac{m(E)c_{j}\eta_{W(E),E,X\setminus(E\cup\{j\})}}{m(X\setminus\{j\})}\quad\\ &-\sum_{\begin{subarray}{c}E\subset X\\ \emptyset\neq C\setminus E\subseteq C^{+}\end{subarray}}(-1)^{\lvert X\setminus E\rvert-1}2^{\lvert E\cup C\rvert-1}\frac{m(E)c_{j}\eta_{W(E),E,X\setminus(E\cup\{j\})}}{m(X\setminus\{j\})}\end{split}$
$=$ $\sum_{F\subseteq X\setminus C}(-1)^{\lvert F\rvert}2^{\lvert X\setminus F\rvert+\ell(F,X\setminus F)}\frac{m(X\setminus F)}{m(X)}\eta_{T,\emptyset,F}r(W(X\setminus F),X\setminus F),$
where $X=F\sqcup(C\cup E)$ and we denote by $r(W(Z),Z)$ the relation (23) for the subset $Z$ of nullity 111 and the connected component $W(Z)$.
This provides an alternative proof of [CDD+20, Theorem 6.13].
@@ 7. Configuration spaces
In this section, we apply the main result of this article to the cohomology ring of ordered configuration spaces in $\mathbb{R}^{b}\times(S^{1})^{a}$.
@@
[…]
@@
Recall that in $H^{*}(G^{n})$, $\psi_{ik}=\psi_{ij}+\psi_{jk}$ and that $\omega_{ij}^{2}=0$ by the product rule. Moreover the relation $\omega_{ij}\omega_{jk}\omega_{ik}=0$ holds multiplying equation (29) (resp. eq. (30)) by $\omega_{ik}$. In the case $b=1$, correction terms appear in our formulas. In the article [CT78] and in this one, the element $\omega_{ij}$ represent the class of the subvariety $H_{ij}$ and hence there is no canonical morphism between the two algebras $\operatorname{gr}_{F_{*}}H^{*}(\operatorname{Conf}_{n}(G;\mathbb{K}))$ and $H^{*}(\operatorname{Conf}_{n}(G);\mathbb{K})$ for $n\geq 3$.